# Estimating non-linear functionals of trawl processes

### v2

@@ Appendix A Computing the AVAR
This case will be slightly different from the others. First observe that in view of
$\sum_{j=l\lor k+1}^{n-1}\mathrm{Leb}(\cup_{i=0}^{(k-l^{\prime})\land(j-l)}\mathcal{P}_{A}^{n}(i,j))\leq\mathrm{Leb}(A),$
we conclude as in previous cases that $\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,1}\rvert^{2})\lesssim n\Delta_{n}$. Second, decompose
$\mathcal{S}_{l,l^{\prime}}^{n,2}=\mathcal{S}_{l,l^{\prime}}^{\star,n}+\tilde{\mathcal{S}}_{l,l^{\prime}}^{n,2},$
where
@@
By arguing exactly in the same way of the proof of Lemma 3 in [17] (specifically the analysis done for $\frac{1}{n\Delta_{n}}\sum_{j=l_{n}}^{n}\mathbb{E}(\xi_{j,n}^{(2)}\xi_{j,n}^{(3)}\mid\mathscr{G}_{j-1}^{n})$ under the notation introduced in that paper) we conclude that $\tilde{\mathcal{S}}_{l,l^{\prime}}^{n,2}$ satisfy (69) and that
$\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{\star,n}\rvert)\lesssim\Delta_{n}(l\lor l^{\prime}).$ (72)
Note that by replacing (69) by (72) in the arguments above, we obtain that $\mathbb{E}(\rvert\mathcal{V}_{n}^{(1,2)}-\mathbb{E}(\mathcal{V}_{n}^{(1,2)})\rvert)\rightarrow 0$, as $n\rightarrow\infty$.
@@
In this case, by applying (71), we deduce that (67) holds with
$\mathcal{S}_{l,l^{\prime}}^{n}$ $=-\sum_{m=0}^{n-1}\sum_{k=l^{\prime}}^{(n-l-1)\land(m+l^{\prime})}\beta_{k,m,l^{\prime}}^{(1)}\theta_{m,k}^{l,l^{\prime}}-\sum_{j=l\lor l^{\prime}+1}^{n-1}\sum_{k=l^{\prime}}^{j-1}\beta_{j,j-1,l}^{(2)}\beta_{k,j-1,l^{\prime}}^{(2)}\mathbb{E}(\alpha_{k+1,j}^{2})$
$=:\mathcal{S}_{l,l^{\prime}}^{n,1}+\mathcal{S}_{l,l^{\prime}}^{n,2}$
where $\theta_{m,k}^{l,l^{\prime}}:=\sum_{j=m\lor(k+l)+1}^{n-1}\mathbb{E}(\alpha_{k+1,j}^{3})$. By arguing exactly in the same way as $\ell=\ell^{\prime}=3$, we deduce that $\mathcal{S}_{l,l^{\prime}}^{n,1}$ satisfies the bound in (69). Now, assume without loss of generality that $l\geq l^{\prime}$ and put
@@
Then, we can further decompose $\mathcal{S}_{l,l^{\prime}}^{n,2}=\sum_{\imath=1}^{3}\mathcal{\tilde{S}}_{l,l^{\prime}}^{n,\imath}$, in which $\mathcal{\tilde{S}}_{l,l^{\prime}}^{n,1}:=-\sum_{j=l\lor l^{\prime}+1}^{n-1}\beta_{j,j-1,l}^{(2)}\vartheta_{j,l,l^{\prime}}^{(1)}$ and
$\mathcal{\tilde{S}}_{l,l^{\prime}}^{n,\imath}:=-\sum_{j=l\lor l^{\prime}+1}^{n-1}[\beta_{j,j-1,l}^{(2)}\vartheta_{j,l,l^{\prime}}^{(\imath)}-\mathbb{E}(\beta_{j,j-1,l}^{(2)}\vartheta_{j,l,l^{\prime}}^{(\imath)})],\,\,\,\imath=2,3.$
Note that $\beta_{j,j-1,l}^{(2)}\vartheta_{j,l,l^{\prime}}^{(\imath)}$ are $l$-dependent. Furthermore, by Rosenthal’s inequality, Lemma 1, and the fact that for $m\leq j-1-l^{\prime}$ it holds that $\sum_{k=i+l^{\prime}}^{m+l^{\prime}}\mathbb{E}(\alpha_{k+1,j}^{2})\lesssim\int_{t_{j}-t_{m+l^{\prime}}}^{t_{j+1}-t_{m+l^{\prime}}}a(s)\mathrm{d}s,$ we infer the following estimate
@@
Likewise, $\mathbb{E}(\rvert\vartheta_{j,l,l^{\prime}}^{(3)}\rvert^{4})\lesssim\Delta_{n}^{4}$. All in all, allow us to conclude as in the case $\ell=\ell^{\prime}=1$ that $\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,2}\rvert^{2})\lesssim n\Delta_{n}(l\Delta_{n})$.
@@
Here we will argue as in the situation when $\ell=1,\ell^{\prime}=2$. We start by noting that (67) is satisfied with
@@
$-\sum_{j=l\lor l^{\prime}}^{n-1}\sum_{k=l^{\prime}}^{j-1}[\beta_{j,j-1,l}^{(2)}\varsigma_{k,j}-\mathbb{E}(\beta_{j,j-1,l}^{(2)}\varsigma_{k,j})]\mathrm{Leb}(\cup_{i=0}^{k-l^{\prime}}\mathcal{P}_{A}^{n}(i,j))$
$=:\mathcal{S}_{l,l^{\prime}}^{n,1}+\mathcal{S}_{l,l^{\prime}}^{n,2}.$
Furthermore, by using that $\sum_{j=l\lor k+1}^{n-1}\mathrm{Leb}(\cup_{i=0}^{(k-l^{\prime})\land(j-l)}\mathcal{P}_{A}^{n}(i,j))\leq\mathrm{Leb}(A)$ and $\mathbb{E}(\varsigma_{k,p}\varsigma_{k,j})\lesssim\Delta_{n}$, we conclude as in the case $\ell=\ell^{\prime}=4$ that $\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,1}\rvert^{2})\lesssim n\Delta_{n}$. Next, we decompose
$\mathcal{S}_{l,l^{\prime}}^{n,2}=\mathcal{S}_{l,l^{\prime}}^{\star,n}+\tilde{\mathcal{S}}_{l,l^{\prime}}^{n,2},$
where
@@
The Cauchy-Swarchz inequality and the fact that $\mathrm{Leb}(\cup_{i=0}^{k-l^{\prime}}\mathcal{P}_{A}^{n}(i,j))=\int_{t_{j}-t_{k-l^{\prime}}}^{t_{j+1}-t_{k-l^{\prime}}}a(s)\mathrm{d}s$ imply that
$\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{\star,n}\rvert)\lesssim\Delta_{n}\sum_{j=l\lor l^{\prime}}^{l+l^{\prime}+1}\left(\sum_{k=l^{\prime}}^{j-1}\int_{t_{j}-t_{k-l^{\prime}}}^{t_{j+1}-t_{k-l^{\prime}}}a(s)\mathrm{d}s\right)^{1/2}\lesssim\Delta_{n}(l\lor l^{\prime}).$
On the other hand, by setting
@@
we can further decompose $\tilde{\mathcal{S}}_{l,l^{\prime}}^{n,2}=\mathcal{\tilde{S}}_{l,l^{\prime}}^{n,1}+\mathcal{\tilde{S}}_{l,l^{\prime}}^{n,2}$, where $\mathcal{\tilde{S}}_{l,l^{\prime}}^{n,1}:=-\sum_{j=l\lor l^{\prime}+1}^{n-1}\beta_{j,j-1,l}^{(2)}\vartheta_{j,l,l^{\prime}}^{(1)}$ and
$\mathcal{\tilde{S}}_{l,l^{\prime}}^{n,2}:=-\sum_{j=l\lor l^{\prime}+1}^{n-1}[\beta_{j,j-1,l}^{(2)}\vartheta_{j,l,l^{\prime}}^{(2)}-\mathbb{E}(\beta_{j,j-1,l}^{(2)}\vartheta_{j,l,l^{\prime}}^{(2)})].$
As in the previous case, $\beta_{j,j-1,l}^{(2)}\vartheta_{j,l,l^{\prime}}^{(\imath)}$ are $l$-dependent and $\mathbb{E}(\rvert\vartheta_{j,l,l^{\prime}}^{(\imath)}\rvert^{4})\lesssim\Delta_{n}^{4}$, reason why $\mathbb{E}(\rvert\tilde{\mathcal{S}}_{l,l^{\prime}}^{n,2}\rvert^{2})\lesssim n\Delta_{n}(l\Delta_{n})$.
@@
Another application of (71) gives us that in this situation that
$\mathcal{S}_{l,l^{\prime}}^{n}=$ $\sum_{k=l^{\prime}+l+1}^{n-1}\chi_{k}\sum_{k^{\prime}=l^{\prime}}^{k-l-1}\beta_{k^{\prime},k-1,l^{\prime}}^{(2)}\theta_{k,k^{\prime}}+\sum_{k=l^{\prime}+l+1}^{n-1}\sum_{k^{\prime}=l^{\prime}}^{k-l-1}[\chi_{k}\beta_{k^{\prime},k,l^{\prime}}^{(1)}-\mathbb{E}(\chi_{k}\beta_{k^{\prime},k,l^{\prime}}^{(1)})]\theta_{k,k^{\prime}}$
$+\sum_{m=l^{\prime}+l+2}^{n-2}\sum_{k=l^{\prime}+l+1}^{m-1}\chi_{k}\sum_{k^{\prime}=l^{\prime}}^{k-l-1}\beta_{k^{\prime},m,l^{\prime}}^{(1)}\theta_{m,k^{\prime}}$
$=:$ $\sum_{\imath=1}^{3}\mathcal{S}_{l,l^{\prime}}^{n,\imath},$
where $\theta_{k,k^{\prime}}:=\sum_{j=k+1}^{n-1}\mathbb{E}[\alpha_{k^{\prime}+1,j}^{2}]$. Once again, each $\mathcal{S}_{l,l^{\prime}}^{n,\imath}$ is a sum of martingale difference. Furthermore, by reasoning as in the proof of Lemma 3 in [17] (see the discussion concerned to $\frac{1}{n\Delta_{n}}\sum_{j=l_{n}}^{n}\mathbb{E}(\xi_{j,n}^{(3)}\xi_{j,n}^{(4)}\mid\mathscr{G}_{j-1}^{n})$), the second moment of each summand in $\mathcal{S}_{l,l^{\prime}}^{n,\imath}$ is bounded up to a constant by $\Delta_{n}$, from which it is clear that $\mathcal{S}_{l,l^{\prime}}^{n}$ satisfies (68).
@@
Obviously in this situation we have
$\mathcal{S}_{l,l^{\prime}}^{n}=-\sum_{j=l\lor l^{\prime}}^{n-1}\sum_{k=l}^{j-1}\sum_{k^{\prime}=l^{\prime}}^{j-1}[\chi_{k}\varsigma_{k^{\prime},j}-\mathbb{E}(\chi_{k}\varsigma_{k^{\prime},j})]\mathrm{Leb}(\cup_{i=0}^{(k^{\prime}-l^{\prime})\land(k-l)}\mathcal{P}_{A}^{n}(i,j)).$
Assume now that $l\geq l^{\prime}$. Then, using that for $k\geq k^{\prime}+1$, it holds
$\chi_{k}=L(\cup_{i=0}^{k^{\prime}}\mathcal{P}_{A}^{n}(i,k))+\sum_{i=k^{\prime}+1}^{k}\alpha_{i,k}=:\lambda_{k^{\prime},k}+\sum_{i=k^{\prime}+1}^{k}\alpha_{i,k}$
we can write $\mathcal{S}_{l,l^{\prime}}^{n}=\sum_{\imath=1}^{4}\mathcal{S}_{l,l^{\prime}}^{n,\imath},$ where we have set
@@
$\mathcal{S}_{l,l^{\prime}}^{n,2}$ $:=-\sum_{j=l}^{n-1}\sum_{k^{\prime}=l}^{j-2}\sum_{i=k^{\prime}+1}^{j-1}\sum_{k=i}^{j-1}[\varsigma_{k^{\prime},j}\alpha_{i,k}-\mathbb{E}(\varsigma_{k^{\prime},j}\alpha_{i,k})]\theta_{k,l,k^{\prime},l^{\prime}}^{j};$
@@
in which $\theta_{k,l,k^{\prime},l^{\prime}}^{j}:=\mathrm{Leb}(\cup_{i=0}^{(k-l)\land(k^{\prime}-l^{\prime})}\mathcal{P}_{A}^{n}(i,j))$ and $\bar{\theta}_{m,k^{\prime},l^{\prime}}:=\mathrm{Leb}(\cup_{i=0}^{(k^{\prime}-l^{\prime})\land(k-l)}\cup_{j=m+1}^{n-1}\mathcal{P}_{A}^{n}(i,j))\leq\mathrm{Leb}(A)$. The same line of arguments used in the case $\ell=\ell^{\prime}=2$ and $\ell=\ell^{\prime}=4$ allow us to conclude that $\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,\imath}\rvert^{2})\lesssim n\Delta_{n}(C_{\alpha}+(n\Delta_{n})^{1-2(\alpha-1)})$, for $\imath=1,2$, respectively. Next, note that $\mathcal{S}_{l,l^{\prime}}^{n,3}$ is once again a sum of martingale differences. Furthermore, for a given fix $k^{\prime}$, $(\alpha_{k^{\prime}+1,m})_{k^{\prime}+1\leq m\leq n-2}$ and $(\lambda_{k^{\prime},k})_{k^{\prime}+1\leq k\leq n-2}$ are independent due to the independent scattered property of $L$. Therefore
@@
After some tedious, yet simple, change of order of summations we get that
@@
Analogous arguments can be used for the case $l<l^{\prime}$.
@@
In this case it is enough to note that by swapping $k$ with $k^{\prime}$ and replacing $\chi_{k}$ by $\varsigma_{k^{\prime},j}$ in the case $\ell=2,\ell^{\prime}=3$, we deduce that (67) holds and $\mathcal{S}_{l,l^{\prime}}^{n}$ satisfies (68).
@@
Let Assumptions of Theorems 5 and 6 hold. Then,
$\Sigma_{n}^{(\ell,\ell^{\prime})}(l,l^{\prime}):=\frac{1}{n\Delta_{n}}\sum_{j=l\lor l^{\prime}}^{n}\mathbb{E}(\xi_{j,l,n}^{(\ell)}\xi_{j,l^{\prime},n}^{(\ell^{\prime})}\mid\mathscr{G}_{j-1}^{n})=F_{n}^{(\ell,\ell^{\prime})}(t_{l},t_{l^{\prime}})+\mathrm{o}(1),\,\,\ell,\ell^{\prime}=1,2,3,4,$ (74)
uniformly on $l,l^{\prime}$ and where $F_{n}^{(\ell,\ell^{\prime})}$ is a sequence of bounded functions such that for all $s,r\geq 0$, $F_{n}^{(\ell,\ell^{\prime})}(\tau_{n}(s),\tau_{n}(r))\rightarrow F^{(\ell,\ell^{\prime})}(s,r)$ as $n\rightarrow\infty$ for some measurable function $F^{(\ell,\ell^{\prime})}$ satisfying that $\sum_{\ell,\ell^{\prime}=1}^{4}F_{n}^{(\ell,\ell^{\prime})}(\tau_{n}(s),\tau_{n}(r))\rightarrow\sum_{\ell,\ell^{\prime}=1}^{4}F^{(\ell,\ell^{\prime})}(s,r)=\Sigma_{a}(s,r),$for all $s,r$.
@@
As in the proof of Lemma 7 we analyze separately the terms $\Sigma_{n}^{(\ell,\ell^{\prime})}(l,l^{\prime})$. Note that in view that $\Sigma_{n}^{(\ell^{\prime},\ell)}(l,l^{\prime})=\Sigma_{n}^{(\ell,\ell^{\prime})}(l^{\prime},l)$, it is enough to show that (74) holds for $\ell=1,2,3,4,$ and $\ell^{\prime}\geq\ell$. Furthermore, it will be implicitly assumed that for fixed $n$, the function $F_{n}^{(\ell,\ell^{\prime})}$ defined below vanishes outside of $0,n\Delta_{n}]$. Finally, unless said otherwise, the error terms (represented as usual by $\mathrm{O}(\cdot)$) presented below will always be uniformly on $l,l^{\prime}$.
@@
$=$ $F_{n,1}^{(\ell,\ell^{\prime})}(t_{l},t_{l^{\prime}})+F_{n,2}^{(\ell,\ell^{\prime})}(t_{l},t_{l^{\prime}})+\mathrm{O}(\Delta_{n}).$
@@
$F_{n,1}^{(1,1)}(s,r)$ $=\mathfrak{K}_{4}\frac{1}{\Delta_{n}}\int_{s\lor r}^{s\lor r+\Delta_{n}}a(u)\mathrm{d}u\left(1-\frac{s\lor r}{n\Delta_{n}}\right)$
$=\mathfrak{K}_{4}a_{n}(s\lor r)\left(1-\frac{s\lor r}{n\Delta_{n}}\right)$

### v3

@@ Appendix A Approximations of the Asymptotic Variance
Since
$\sum_{j=l\lor k+1}^{n-1}\mathrm{Leb}(\cup_{i=0}^{(k-l^{\prime})\land(j-l)}\mathcal{P}_{A}^{n}(i,j))\leq\mathrm{Leb}(A),$ (72)
we conclude, as in previous cases, that $\mathcal{S}_{l,l^{\prime}}^{n,1}$ satisfies (68). The second term requires a separate argument. Decompose $\mathcal{S}_{l,l^{\prime}}^{n,2}=\mathcal{S}_{l,l^{\prime}}^{\star,n}+\tilde{\mathcal{S}}_{l,l^{\prime}}^{n,2}$, where
@@
Arguing as in the proof of Lemma 3 in [19] (specifically the analysis done for $\mathbb{E}(\xi_{j,n}^{(2)}\xi_{j,n}^{(3)}\mid\mathscr{G}_{j-1}^{n})$ under the notation introduced in that paper) we conclude that $\tilde{\mathcal{S}}_{l,l^{\prime}}^{n,2}$ satisfy (69) and that
$\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{\star,n}\rvert)\lesssim\Delta_{n}((l\lor l^{\prime})+1).$ (73)
With (73) in place of (69), the same argument yields $\mathbb{E}(|\mathcal{V}_{n}^{(1,2)}-\mathbb{E}(\mathcal{V}_{n}^{(1,2)})|)\to 0$, as $n\to\infty$.
@@
By applying (70), we obtain that (67) holds with
$\mathcal{S}_{l,l^{\prime}}^{n}=$ $-\sum_{m=0}^{n-1}\sum_{k=l^{\prime}}^{(n-l-1)\land(m+l^{\prime})}\beta_{k,m,l^{\prime}}^{(1)}\theta_{m,k}^{l,l^{\prime}}$
$-\sum_{j=l\lor(l^{\prime}+1)}^{n-1}\sum_{k=l^{\prime}}^{j-1}[\beta_{j,j-1,l}^{(2)}\beta_{k,j-1,l^{\prime}}^{(2)}-\mathbb{E}(\beta_{j,j-1,l}^{(2)}\beta_{k,j-1,l^{\prime}}^{(2)})]\mathbb{E}(\alpha_{k+1,j}^{2})$
$=:\mathcal{S}_{l,l^{\prime}}^{n,1}+\mathcal{S}_{l,l^{\prime}}^{n,2},$
where $\theta_{m,k}^{l,l^{\prime}}:=\sum_{j=m\lor(k+l)+1}^{n-1}\mathbb{E}(\alpha_{k+1,j}^{3})$. As in the case $\ell=\ell^{\prime}=3$, one deduces that $\mathcal{S}_{l,l^{\prime}}^{n,1}$ satisfies (69). Assume without loss of generality that $l\geq l^{\prime}$, and put
@@
Then $\mathcal{S}_{l,l^{\prime}}^{n,2}=\sum_{\imath=1}^{3}\mathcal{\tilde{S}}_{l,l^{\prime}}^{n,\imath}$, where $\mathcal{\tilde{S}}_{l,l^{\prime}}^{n,1}:=-\sum_{j=l\lor l^{\prime}+1}^{n-1}\beta_{j,j-1,l}^{(2)}\vartheta_{j,l,l^{\prime}}^{(1)}$, and for $\imath=2,3$
$\mathcal{\tilde{S}}_{l,l^{\prime}}^{n,\imath}:=-\sum_{j=l\lor l^{\prime}+1}^{n-1}[\beta_{j,j-1,l}^{(2)}\vartheta_{j,l,l^{\prime}}^{(\imath)}-\mathbb{E}(\beta_{j,j-1,l}^{(2)}\vartheta_{j,l,l^{\prime}}^{(\imath)})].$
Note that $\beta_{j,j-1,l}^{(2)}\vartheta_{j,l,l^{\prime}}^{(\imath)}$ are $l$-uncorrelated. Furthermore, by Rosenthal’s inequality, Lemma 1, and the bound
$\sum_{k=i+l^{\prime}}^{m+l^{\prime}}\mathbb{E}(\alpha_{k+1,j}^{2})\lesssim\int_{t_{j}-t_{m+l^{\prime}}}^{t_{j+1}-t_{m+l^{\prime}}}a(s)\mathrm{d}s,\quad m\leq j-1-l^{\prime},$
we obtain
@@
Likewise, $\mathbb{E}(\rvert\vartheta_{j,l,l^{\prime}}^{(3)}\rvert^{4})\lesssim\Delta_{n}^{4}$. Consequently, arguing as in the case $\ell=\ell^{\prime}=1$, we conclude that $\mathcal{S}_{l,l^{\prime}}^{n,2}$ fulfils (69).
@@
We start by noting that (67) is satisfied with
@@
$-\sum_{j=l\lor l^{\prime}}^{n-1}\sum_{k=l^{\prime}}^{j-1}[\beta_{j,j-1,l}^{(2)}\varsigma_{k,j}-\mathbb{E}(\beta_{j,j-1,l}^{(2)}\varsigma_{k,j})]\mathrm{Leb}(\cup_{i=0}^{k-l^{\prime}}\mathcal{P}_{A}^{n}(i,j))=:\mathcal{S}_{l,l^{\prime}}^{n,1}+\mathcal{S}_{l,l^{\prime}}^{n,2}.$
Using (72) and $\mathbb{E}(\varsigma_{k,p}\varsigma_{k,j})\lesssim\Delta_{n}$, it follows that $\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,1}\rvert^{2})\lesssim n\Delta_{n}$. Next, decompose $\mathcal{S}_{l,l^{\prime}}^{n,2}=\mathcal{S}_{l,l^{\prime}}^{\star,n}+\mathcal{S}_{l,l^{\prime}}^{\star\star,n}$, where
@@
By setting
@@
we can further decompose $\mathcal{S}_{l,l^{\prime}}^{\star\star,n}=\mathcal{\tilde{S}}_{l,l^{\prime}}^{n,1}+\mathcal{\tilde{S}}_{l,l^{\prime}}^{n,2}$, where
$\mathcal{\tilde{S}}_{l,l^{\prime}}^{n,1}:=-\sum_{j=l+l^{\prime}+2}^{n-1}\beta_{j,j-1,l}^{(2)}\vartheta_{j,l,l^{\prime}}^{(1)},$
and
$\mathcal{\tilde{S}}_{l,l^{\prime}}^{n,2}:=-\sum_{j=l+l^{\prime}+2}^{n-1}[\beta_{j,j-1,l}^{(2)}\vartheta_{j,l,l^{\prime}}^{(2)}-\mathbb{E}(\beta_{j,j-1,l}^{(2)}\vartheta_{j,l,l^{\prime}}^{(2)})].$
As in the previous case, $\beta_{j,j-1,l}^{(2)}\vartheta_{j,l,l^{\prime}}^{(\imath)}$ are $(l+1)$-correlated and $\mathbb{E}(\rvert\vartheta_{j,l,l^{\prime}}^{(\imath)}\rvert^{4})\lesssim\Delta_{n}^{4}$, reason why $\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{\star\star,n}\rvert^{2})\lesssim n\Delta_{n}((l+1)\Delta_{n})$. It remains to treat $\mathcal{S}_{l,l^{\prime}}^{\star,n}$. By the Cauchy-Swarchz inequality and the fact that $\mathrm{Leb}(\cup_{i=0}^{k-l^{\prime}}\mathcal{P}_{A}^{n}(i,j))=\int_{t_{j}-t_{k-l^{\prime}}}^{t_{j+1}-t_{k-l^{\prime}}}a(s)\mathrm{d}s$ we deduce that
$\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{\star,n}\rvert)\lesssim\Delta_{n}\sum_{j=l\lor l^{\prime}}^{l+l^{\prime}+1}\left(\sum_{k=l^{\prime}}^{j-1}\int_{t_{j}-t_{k-l^{\prime}}}^{t_{j+1}-t_{k-l^{\prime}}}a(s)\mathrm{d}s\right)^{1/2}\lesssim\Delta_{n}((l\lor l^{\prime})+1).$
Thus, $\mathcal{S}_{l,l^{\prime}}^{\star,n}$ satisfies (73), which, as explained above, suffices.
@@
Another application of (70) yields
$\mathcal{S}_{l,l^{\prime}}^{n}=$ $\sum_{k=l^{\prime}+l+1}^{n-2}\chi_{k}\sum_{k^{\prime}=l^{\prime}}^{k-l-1}\beta_{k^{\prime},k-1,l^{\prime}}^{(2)}\theta_{k,k^{\prime}}+\sum_{k=l^{\prime}+l+1}^{n-2}\sum_{k^{\prime}=l^{\prime}}^{k-l-1}[\chi_{k}\beta_{k^{\prime},k,l^{\prime}}^{(1)}-\mathbb{E}(\chi_{k}\beta_{k^{\prime},k,l^{\prime}}^{(1)})]\theta_{k,k^{\prime}}$
$+\sum_{m=l^{\prime}+l+2}^{n-2}\sum_{k=l^{\prime}+l+1}^{m-1}\chi_{k}\sum_{k^{\prime}=l^{\prime}}^{k-l-1}\beta_{k^{\prime},m,l^{\prime}}^{(1)}\theta_{m,k^{\prime}}=:\sum_{\imath=1}^{3}\mathcal{S}_{l,l^{\prime}}^{n,\imath},$
where $\theta_{k,k^{\prime}}:=\sum_{j=k+1}^{n-1}\mathbb{E}[\alpha_{k^{\prime}+1,j}^{2}]$. Once again, each $\mathcal{S}_{l,l^{\prime}}^{n,\imath}$ is a sum of martingale differences. By reasoning as in the proof of Lemma 3 in [19] (see the treatment of $\frac{1}{n\Delta_{n}}\sum_{j=l_{n}}^{n}\mathbb{E}(\xi_{j,n}^{(3)}\xi_{j,n}^{(4)}\mid\mathscr{G}_{j-1}^{n})$), the second moment of each summand in $\mathcal{S}_{l,l^{\prime}}^{n,\imath}$ is $\mathrm{O}(\Delta_{n})$. Hence, $\mathcal{S}_{l,l^{\prime}}^{n,\imath}$ satisfies (68), for every $\imath=1,2,3$.
@@
Using (71), we have
$\mathcal{S}_{l,l^{\prime}}^{n}=-\sum_{j=l\lor l^{\prime}}^{n-1}\sum_{k=l}^{j-1}\sum_{k^{\prime}=l^{\prime}}^{j-1}[\chi_{k}\varsigma_{k^{\prime},j}-\mathbb{E}(\chi_{k}\varsigma_{k^{\prime},j})]\theta_{k,l,k^{\prime},l^{\prime}}^{j}.$
Assume $l\geq l^{\prime}$ and note that for $k\geq k^{\prime}+1$, it holds
$\chi_{k}=L(\cup_{i=0}^{k^{\prime}}\mathcal{P}_{A}^{n}(i,k))+\sum_{i=k^{\prime}+1}^{k}\alpha_{i,k}=:\lambda_{k^{\prime},k}+\sum_{i=k^{\prime}+1}^{k}\alpha_{i,k}.$
Then $\mathcal{S}_{l,l^{\prime}}^{n}=\sum_{\imath=1}^{4}\mathcal{S}_{l,l^{\prime}}^{n,\imath}$, where
@@
$\mathcal{S}_{l,l^{\prime}}^{n,2}$ $:=-\sum_{i=l+1}^{n-2}\sum_{k^{\prime}=l}^{i-1}\sum_{j=i}^{n-1}\sum_{k=i}^{j-1}[\alpha_{i,k}\varsigma_{k^{\prime},j}-\mathbb{E}(\varsigma_{k^{\prime},j}\alpha_{i,k})]\theta_{k,l,k^{\prime},l^{\prime}}^{j};$
@@
with
$\bar{\theta}_{m,k,l,k^{\prime},l^{\prime}}:=\mathrm{Leb}(\cup_{i=0}^{(k^{\prime}-l^{\prime})\land(k-l)}\cup_{j=m+1}^{n-1}\mathcal{P}_{A}^{n}(i,j))\leq\mathrm{Leb}(A).$
The same line of arguments used in the cases $\ell=\ell^{\prime}=2$ and $\ell=\ell^{\prime}=4$, result in $\mathbb{E}(\rvert\mathcal{S}_{l,l^{\prime}}^{n,\imath}\rvert^{2})\lesssim n\Delta_{n}(C_{\alpha}+(n\Delta_{n})^{1-2(\alpha-1)})$, for $\imath=1,2$. Next, note that $\mathcal{S}_{l,l^{\prime}}^{n,3}$ is once again a sum of martingale differences. For $\imath=3$, using independence of $(\alpha_{k^{\prime}+1,m})_{k^{\prime}+1\leq m\leq n-2}$ and $(\lambda_{k^{\prime},k})_{k^{\prime}+1\leq k\leq n-2}$ for fixed $k^{\prime}$, we deduce
@@
After changing the order of summations, we get that
@@
The case $l<l^{\prime}$ follows analogously.
@@
This case follows from $(\ell,\ell^{\prime})=(2,3)$ by swapping $k$ and $k^{\prime}$ and replacing $\chi_{k}$ with $\varsigma_{k^{\prime},j}$. Hence, (67) holds and each $\mathcal{S}_{l,l^{\prime}}^{n,\iota}$ satisfies (68).
@@
Let Assumptions of Theorems 5 and 6 hold. Then, for all $\ell,\ell^{\prime}=1,2,3,4$
$\Sigma_{n}^{(\ell,\ell^{\prime})}(l,l^{\prime}):=\frac{1}{n\Delta_{n}}\sum_{j=l\lor l^{\prime}}^{n}\mathbb{E}(\xi_{j,l,n}^{(\ell)}\xi_{j,l^{\prime},n}^{(\ell^{\prime})})=F_{n}^{(\ell,\ell^{\prime})}(t_{l},t_{l^{\prime}})+\mathrm{o}(1),\,\,$ (75)
uniformly on $l,l^{\prime}$, where $F_{n}^{(\ell,\ell^{\prime})}$ is a sequence of bounded functions such that for all $s,r\geq 0$, $F_{n}^{(\ell,\ell^{\prime})}(\tau_{n}(s),\tau_{n}(r))\rightarrow F^{(\ell,\ell^{\prime})}(s,r)$ as $n\rightarrow\infty$. Moreover,
$\sum_{\ell,\ell^{\prime}=1}^{4}\Sigma_{n}^{(\ell,\ell^{\prime})}(\tau_{n}(s),\tau_{n}(r))\rightarrow\Sigma_{a}(s,r),\,\,\forall\,s,r\geq 0,$ (76)
where $\Sigma_{a}$ as in (14).
@@
As in the proof of Lemma 7, we analyse $\Sigma_{n}^{(\ell,\ell^{\prime})}(l,l^{\prime})$ case by case. In view of $\Sigma_{n}^{(\ell^{\prime},\ell)}(l,l^{\prime})=\Sigma_{n}^{(\ell,\ell^{\prime})}(l^{\prime},l)$, it is enough to consider $\ell^{\prime}\geq\ell$. Furthermore, for each pair $(\ell,\ell^{\prime})$, we construct a function $F_{n}^{(\ell,\ell^{\prime})}$ such that (75) holds. We extend $F_{n}^{(\ell,\ell^{\prime})}$ by zero outside $0,n\Delta_{n}]$. Unless stated otherwise, all error terms $\mathrm{O}(\cdot)$ are uniform in $l,l^{\prime}$.
@@
Thus,
$\Sigma_{n}^{(1,1)}(l,l^{\prime})=F_{n,1}^{(1,1)}(t_{l},t_{l^{\prime}})+F_{n,2}^{(1,1)}(t_{l},t_{l^{\prime}})+\mathrm{O}(\Delta_{n}),$
@@
$F_{n,1}^{(1,1)}(s,r)=\mathfrak{K}_{4}a_{n}(s\lor r)\Big(1-\frac{s\lor r}{n\Delta_{n}}\Big),$